OpenStudy (anonymous):

Factor the Following Polynomials: x²-81 x²-13x+36 Which polynomial is a special product and why? I dont know how to factor. Anyone want to explain?

OpenStudy (anonymous):

in english, this is the "difference to two squares"

OpenStudy (anonymous):

the difference of two squares factors as \[a^2-b^2=(a+b)(a-b)\]

OpenStudy (anonymous):

so would i plug into that ^^

OpenStudy (anonymous):

so the first example, we have \(x^2-81\) think \(a=x,b=9\)

OpenStudy (anonymous):

yes, replace \(a\) by \(x\) and \(b\) by \(9\)

OpenStudy (anonymous):

\[x^2-13x+36\] is a bit different you have to think of two numbers whose product is \(36\) and that add to \(-13\) since they add to \(-13\) they must both be negative

OpenStudy (anonymous):

x^2-9^2= (x+9 ) (x-9)

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

now for \(x^2-13x+36\) since \(-9\times (-3)=36\) and \(-9+(-3)=-13\) it factors as \((x-9)(x-3)\) you should check that this works